Just as a car start to accelerate from rest at a constant rate of 2.44m/s2, a bus moving at a constant speed of 19.6m/s passes a car in a parallel lane.

(a) How long before the car overtakes the bus?

(b) How fast is the car going then?

(c) How far has the car gone at that point?



1
Expert's answer
2021-02-28T07:27:19-0500

(a) The distance traveled by the car can be written as follows:


"d_1=\dfrac{1}{2}at^2."

The distance traveled by the bus can be written as follows:


"d_2=vt."

When the car overtakes the bus "d_1=d_2" and we can write:


"\dfrac{1}{2}at^2=vt,""t=\dfrac{2v}{a}=\dfrac{2\cdot19.6\ \dfrac{m}{s}}{2.44\ \dfrac{m}{s^2}}=16.1\ s."

(b) We can find the speed of the car when the bus overtaking from the kinematic equation:


"v=v_0+at=0+2.44\ \dfrac{m}{s^2}\cdot16.1\ s=39.3\ \dfrac{m}{s}."

(c) We can find the distance that the car travels at that point from the kinematic equation:


"d_1=\dfrac{1}{2}at^2,""d_1=\dfrac{1}{2}\cdot2.44\ \dfrac{m}{s^2}\cdot(16.1\ s)^2=316\ m."